ANALYSIS OF THE HIGH-RESOLUTION JET-COOLED FLUORESCENCE EXCITATION SPECTRA OF THE HIGHER LEVELS IN THE 29 769 CM^{−1} BAND SYSTEM OF ACETALDEHYDE: FURTHER REFINE-MENTS OF THE TORSION-INVERSION POTENTIAL CONSTANTS.

H. LIU, E. C. LIM, Department of Chemistry, University of Akron, Akron, Ohio 44325-3601; C. MUNOZ-CARO, A. NINO, E. U. Informática, Universidad de Castilla-La Mancha, Ronda de Calatrava 5, 13071 Ciudad Real, Spain; R. H. JUDGE, Department of Chemistry, University of Wisconsin-Parkside, Kenosha, WI 53141-2000; \underline{D. C. MOULE} , Department of Chemistry, Brock University, St. Catharines, Ontario, Canada, L2S 3A1.

Acetaldehyde molecules in the first singlet excited state contain two coupled low frequency vibrational modes, ν_{14} (inversion) and ν_{15} (torsion). In an earlier paper, ^{a} a two dimensional least squares analysis of the first seven bands in the UV spectrum of acetaldehyde jet-cooled vapor gave an excited state barrier to inversion of 585.13 cm^{−1} and a torsional barrier of 721.43 cm^{−1}. A discussion of the results of a vibrational and rotational analysis of 12 additional bands beyond this earlier analysis will be presented. The assignment of the 14_{0}^{2} band has changed and the inclusion of the higher members of the torsional mode has given refined values for the barriers.

aH. Liu, E. C. Lim, C. Muñoz-Caro, A. Niño, R. H. .Judge and D. C. Moule, J. Mol. Spectrosc., 175, 172-189 (1996).